OpenStudy (anonymous):

a perfume sample in a small cylindrical bottle which is 0.8 cm in diameter is distributed for free. How tall is the bottle if it contains 2 mL of the perfume

OpenStudy (anonymous):

*Note that: 2mL is equivalent to 2 cubic centimeters. Volume of bottle = 2 cm^3 Volume is given by, V = circular area * height Circular area is given by, A = pi * r^2, where r is the radius. Calculate the circular area, A= pi * (0.4)^2 (note that 0.8 is the diameter)( half of that is the radius) A= 0.16pi Volume = 2 = 0.16pi *height Height = 2/(0.16pi) = 3.978cm = 3.98cm approx.